OpenStudy (anonymous):

what is the number of sides of a polygon whose measure of an interior angle is 156?

OpenStudy (anonymous):

15 sides.

OpenStudy (anonymous):

If each side of the polygon is 156 degrees, set 156 = (n-2) × 180° / n and solve for n, where n is the number of sides.

OpenStudy (jdoe0001):

\(\large \begin{array}{rllll} &&\textit{number of sides}\\ &\downarrow&\downarrow\\ 156\cdot &n=(&n-2)\cdot 180\\ \uparrow\\ \textit{sum of interior angles} \end{array}\) solve for "n"